The earth is on average 150 million km from the sun calculate it's average speed in orbit
1
Expert's answer
2021-10-20T10:41:51-0400
"v=L/T=2\pi R/T,\\
v=2\pi·(150·10^6)/(365·24·3600)=30\text{ km/s}."